OpenStudy (anonymous):

can anyone walk me through this problem: 5x-1+10x^2+7x-2x-4 ?

OpenStudy (anonymous):

We can't solve it without a =somehwere

OpenStudy (anonymous):

Or do we need to simplify it?

OpenStudy (anonymous):

yea simplify

OpenStudy (anonymous):

5x-1+10x^2+7x-2x-4 = 10x^2 +5x +7x -2x - 1 - 4

OpenStudy (anonymous):

by simply rearranging it

OpenStudy (anonymous):

It is simple just collect like terms

OpenStudy (anonymous):

ok thanks guyssss

OpenStudy (anonymous):

\[5x-1+10x^2+7x-2x-4\] \(x^2\) has no match \[7x+5x-2x=10x\] and \[-1-4=-5\] giving you \[10x^2+10x-5\]